How many electrons per second pass through a section of wire carrying a current of 0.70 A.?

What is the current through an 8.0 omega toaster when it is operating on 120 V?

An electron gun in a tv set shoots out a beam of electrons. The beam current is 1.0 × 10 raised to -5. A. HOW MANY ELECTRONS STRIKE THE TV EACH SECOND ? B. HOW MUCH CHARGE STRIKES THE SCREEN IN A MINUTE ?
1
Expert's answer
2019-02-18T14:52:08-0500

The number of electrons per second

N=q/e=It/e=(0.70A×1s)/(1.6×〖10〗^(-19) C)=4.38×〖10〗^18

The Ohm’s law states

I=V/R=120V/8.0Ω=15A

The number of electrons per second

N=q/e=It/e=(1.0×〖10〗^(-5) A×1s)/(1.6×〖10〗^(-19) C)=6.25×〖10〗^13

The total charge

q=It=1.0×〖10〗^(-5) A×60s=6.0×〖10〗^(-4) C=60 mC